PLEASE !!!!!!!!!!!HELP !!!!!!!!!!!!!!!1
If a+b+c+d=45 and a+2=b-2=2c=d/2
FIND the values of a b c d PLEASEEEEEEEEEEEEEEEEE !!!!!!!!!! GOD BLESS YOU !!!!!!!

Answer by ptaylor(2198)

If a+b+c+d=45 and a+2=b-2=2c=d/2
FIND the values of a b c d PLEASEEEEEEEEEEEEEEEEE !!!!!!!!!! GOD BLESS YOU !!!!!!!
a+2=b-2 add 2 to each side
a+2+2=b-2+2 collect like terms
b=a+4
a+2=2c divide each side by 2
c=(a+2)/2
a+2=d/2 multiply each side by 2
2(a+2)=d or
d=2(a+2)
Substituting the above values for b, c, d into the original equation:
a+b+c+d=45 and we get:
a+a+4+(a+2)/2+2(a+2)=45 multiply each term by 2
2a+2a+8+a+2+4a+8=90 subtract 18 from each side
9a=90-18
9a=72 divide both sides by 9
a=8
b=a+4=8+4=12
c=(a+2)/2=(8+2)/2=10/2=5
d=2(a+2)=2(8+2)=2*10=20
CK
a+b+c+d=45
8+12+5+20=45
45=45
and
a+2=b-2
8+2=12-2
10=10
and
a+2=2c
8+2=2*5
10=10
and
a+2=d/2
8+2=20/2
10=10

Hope this helps---ptaylor